# KAN: Kolmogorov–Arnold Networks

### v2

@@ 2.5.1 Simplification techniques
2. Visualization. When we visualize a KAN, to get a sense of magnitudes, we set the transparency of an activation function $\phi_{l,i,j}$ proportional to ${\rm tanh}(\beta A_{l,i,j})$ where $\beta=3$ Hence, functions with small magnitude appear faded out to allow us to focus on important ones.
@@
[…]
@@ 3.1 Toy datasets
We train these KANs by increasing grid points every 200 steps, in total covering $G=\{3,5,10,20,50,100,200,500,1000\}$. We train MLPs with different depths and widths as baselines. Both MLPs and KANs are trained with LBFGS for 1800 steps in total. We plot test RMSE as a function of the number of parameters for KANs and MLPs in Figure 3.1, showing that KANs have better scaling curves than MLPs, especially for the high-dimensional example. For comparison, we plot the lines predicted from our KAN theory as red dashed ($\alpha=k+1=4$), and the lines predicted from Sharma & Kaplan [17] as black-dashed ($\alpha=(k+1)/d=4/d$). KANs can almost saturate the steeper red lines, while MLPs struggle to converge even as fast as the slower black lines and plateau quickly. We also note that for the last example, the 2-Layer KAN $[4,9,1]$ behaves much worse than the 3-Layer KAN (shape $[4,2,2,1]$). This highlights the greater expressive power of deeper KANs, which is the same for MLPs: deeper MLPs have more expressive power than shallower ones.
@@ 3.2 Special functions
[…]
@@
Name scipy.special API Minimal KAN shape test RMSE $<10^{-2}$ Minimal KAN test RMSE Best KAN shape Best KAN test RMSE MLP test RMSE
@@ 3.3 Feynman datasets
I.15.3x $\frac{x-ut}{\sqrt{1-(\frac{u}{c})^{2}}}$ $\frac{1-a}{\sqrt{1-b^{2}}}$ $a,b$ [2,2,1,1] [2,1,1] [2,2,1,1,1] $7.35\times 10^{-3}$ $1.58\times 10^{-3}$ $1.14\times 10^{-3}$ ${\bf 8.54\times 10^{-4}}$
I.16.6 $\frac{u+v}{1+\frac{uv}{c^{2}}}$ $\frac{a+b}{1+ab}$ $a,b$ [2,2,2,2,2,1] [2,2,1] [2,2,1] $1.06\times 10^{-3}$ $1.19\times 10^{-3}$ $1.53\times 10^{-3}$ ${\bf 6.20\times 10^{-4}}$
@@
II.2.42 $\frac{k(T_{2}-T_{1})A}{d}$ $(a-1)b$ $a,b$ [2,2,1] [2,2,1] [2,2,2,1] $8.54\times 10^{-4}$ $7.22\times 10^{-4}$ $1.22\times 10^{-3}$ ${\bf 1.81\times 10^{-4}}$
@@
[…]
@@
    MLPs with fixed width 20, depths swept in $\{2,3,4,5,6\}$, and activations chosen from $\{{\rm Tanh},{\rm ReLU},{\rm SiLU}\}$.
@@ 3.4 Solving partial differential equations
We compare the KAN architecture with that of MLPs using the same hyperparameters $n_{i}=10000$, $n_{b}=800$, and $\alpha=0.01$. We measure both the error in the $L^{2}$ norm and energy ($H^{1}$) norm and see that KAN achieves a much better scaling law with a smaller error, using smaller networks and fewer parameters; see Figure 3.3. Therefore we speculate that KANs might have the potential of serving as a good neural network representation for model reduction of PDEs.
@@ 3.5 Continual Learning
Here we simply present our preliminary results on an extremely simple example, to demonstrate how one could possibly leverage locality in KANs (thanks to spline parametrizations) to reduce catastrophic forgetting. However, it remains unclear whether our method can generalize to more realistic setups, which we leave for future work. We would also like to study how our method can be connected to and combined with SOTA methods in continual learning [27, 28].
@@ 4.2 Unsupervised toy dataset
If an algorithm for solving the unsupervised problem can be devised, it has a considerable advantage over the supervised problem, since it requires only the sets of features $S=(x_{1},x_{2},\cdots,x_{d})$. The supervised problem, on the other hand, tries to predict subsets of features in terms of the others, i.e. it splits $S=S_{\text{in}}\cup S_{\text{out}}$ into input and output features of the function to be learned. Without domain expertise to advise the splitting, there are $22^{d}-$ 2 possibilities such that $|S_{\text{in}}|>0$ and $|S_{\text{out}}|>0$. This exponentially large space of supervised problems can be avoided by using the unsupervised approach. This unsupervised learning approach will be valuable to the knot dataset in Section 4.3. A Google Deepmind team [29] manually chose signature to be the target variable, otherwise they would face this combinatorial problem described above. This raises the question whether we can instead tackle the unsupervised learning directly. We present our method and a toy example below.
We tackle the unsupervised learning problem by turning it into a supervised learning problem on all of the $d$ features, without requiring the choice of a splitting. The essential idea is to learn a function $f(x_{1},\dots,x_{d})=0$ such that $f$ is not the 00-function. To do this, similar to contrastive learning, we define positive samples and negative samples: positive samples are feature vectors of real data. Negative samples are constructed by feature corruption. To ensure that the overall feature distribution for each topological invariant stays the same, we perform feature corruption by random permutation of each feature across the entire training set. Now we want to train a network $g$ such that $g(\mathbf{x}_{\rm real})=1$ and $g(\mathbf{x}_{\rm fake})=0$ which turns the problem into a supervised problem. However, remember that we originally want $f(\mathbf{x}_{\rm real})=0$ and $f(\mathbf{x}_{\rm fake})\neq 0$. We can achieve this by having $g=\sigma\circ f$ where $\sigma(x)={\rm exp}(-\frac{x^{2}}{2w^{2}})$ is a Gaussian function with a small width $w$, which can be conveniently realized by a KAN with shape $[...,1,1]$ whose last activation is set to be the Gaussian function $\sigma$ and all previous layers form $f$. Except for the modifications mentioned above, everything else is the same for supervised training.
@@
Now we demonstrate that the unsupervised paradigm works for a synthetic example. Let us consider a 6D dataset, where $(x_{1},x_{2},x_{3})$ are dependent variables such that $x_{3}=\exp(\sin(x_{1})+x_{2}^{2})$; $(x_{4},x_{5})$ are dependent variables with $x_{5}=x_{4}^{3}$; $x_{6}$ is independent of the other variables. In Figure 4.2, we show that for seed = 0, KAN reveals the functional dependence among $x_{1}$,$x_{2}$, and $x_{3}$; for another seed = 2024, KAN reveals the functional dependence between $x_{4}$ and $x_{5}$. Our preliminary results rely on randomness (different seeds) to discover different relations; in the future we would like to investigate a more systematic and more controlled way to discover a complete set of relations. Even so, our tool in its current status can provide insights for scientific tasks. We present our results with the knot dataset in Section 4.3.
@@ 4.3 Application to Mathematics: Knot Theory
Knot theory is a subject in low-dimensional topology that sheds light on topological aspects of three-manifolds and four-manifolds and has a variety of applications, including in biology and topological quantum computing. Mathematically, a knot $K$ is an embedding of $S^{1}$ into $S^{3}$. Two knots $K$ and $K^{\prime}$ are topologically equivalent if one can be deformed into the other via deformation of the ambient space $S^{3}$, in which case we write $[K]=[K^{\prime}]$. Some knots are topologically trivial, meaning that they can be smoothly deformed to a standard circle. Knots have a variety of deformation-invariant features $f$ called topological invariants, which may be used to show that two knots are topologically inequivalent, $[K]\neq[K^{\prime}]$ if $f(K)\neq f(K^{\prime})$. In some cases the topological invariants are geometric in nature. For instance, a hyperbolic knot $K$ has a knot complement $S^{3}\setminus K$ that admits a canonical hyperbolic metric $g$ such that $\text{vol}_{g}(K)$ is a topological invariant known as the hyperbolic volume. Other topological invariants are algebraic in nature, such as the Jones polynomial.
@@
Table 4: KANs can achieve better accuracy than MLPs with much fewer parameters in the signature classification problem.
@@
D $-0.09+1.04{\rm exp}(-9.59(-0.62{\rm sin}(0.61\mu_{r}+7.26))-0.32{\rm tan}(0.03\lambda-6.59)+1-0.11e^{-1.77(0.31-\mu_{i})^{2})^{2}}-1.09e^{-7.6(0.65(1-0.01\lambda)^{3}}+0.27{\rm atan}(0.53\mu_{i}-0.6)+0.09+{\rm exp}(-2.58(1-0.36\mu_{r})^{2}))$ $[3,2,1]$ KAN 84.0% 0.947 0.997
@@
    The second group of variables involve cusp volume $V$, real part of meridinal translation $\mu_{r}$ and longitudinal translation $\lambda$. Their activations all look like logarithmic functions (which can be verified by the implied symbolic functionality in Section 2.5.1). So the relation is $-\log V+\log\mu_{r}+\log\lambda=0$ which is equivalent to $V=\mu_{r}\lambda$, which is true by definition. It is, however, reassuring that we discover this relation without any prior knowledge.
@@ 4.4 Application to Physics: Anderson localization
[…]
@@
KAN auto 13.99sin(0.28sin($013.99{\rm sin}(0.28{\rm sin}(0.87\lambda+2.22)-0.84{\rm arctan}(0.58E-0.26)+0.85{\rm arctan}(0.94p+0.13)-8.14)-16.74+43.08{\rm exp}(-0.93(0.06(0.13-p)^{2}-0.27{\rm tanh}(0.65E+0.25)+0.63{\rm arctan}(0.54\lambda-0.62)+1)^{2})=$ 0 97.1%
KAN man (step 2) + auto 4.19($04.19(0.28{\rm sin}(0.97\lambda+2.17)-0.77{\rm arctan}(0.83E-0.19)+{\rm arctan}(0.97p+0.15)-0.35)^{2}-28.93+39.27{\rm exp}(-0.6(0.28{\rm cosh}^{2}(0.49p-0.16)-0.34{\rm arctan}(0.65E+0.51)+0.83{\rm arctan}(0.54\lambda-0.62)+1)^{2})=$ 0 97.7%
KAN man (step 3) + auto $-4.63E-10.25(-0.94{\rm sin}(0.97\lambda-6.81)+{\rm tanh}(0.8p-0.45)+0.09)^{2}+11.78{\rm sin}(0.76p-1.41)+22.49{\rm arctan}(1.08\lambda-1.32)+31.72=0$ 97.7%
KAN man (step 4A) $06.92E-6.23(-0.92\lambda-1)^{2}+2572.45(-0.05\lambda+0.95{\rm cosh}(0.11p+0.4)-1)^{2}-12.96{\rm cosh}^{2}(0.53p+0.16)+19.89=$ 0 96.6%
KAN man (step 4B) $07.25E-8.81(-0.83\lambda-1)^{2}-4.08(-p-0.04)^{2}+12.71(-0.71\lambda+(0.3p+1)^{2}-0.86)^{2}+10.29=$ 0 95.4%
@@
Let us assume Alice wants to figure out the mobility edge for MAAM. This task is more complicated and requires more human wisdom. As in the last example, Alice starts from a $[4,2,1,1]$ KAN and trains it but gets an accuracy around 75% which is less than acceptable. She then chooses a larger $[4,3,1,1]$ KAN and successfully gets 98.4% which is acceptable (Figure 4.6 (b) step 1). Alice notices that $\phi$ is not picked up on by KANs, which means that the mobility edge is independent of the phase factor $\phi$ (agreeing with Eq. (4.11)). If Alice turns on the automatic symbolic regression (using a large library consisting of exp, tanh etc.), she would get a complicated formula in Tabel 6-MAAM-KAN auto, which has 97.1% accuracy. However, if Alice wants to find a simpler symbolic formula, she will want to use the manual mode where she does the symbolic snapping by herself. Before that she finds that the $[4,3,1,1]$ KAN after training can then be pruned to be $[4,2,1,1]$, while maintaining $97.7\%$ accuracy (Figure 4.6 (b)). Alice may think that all activation functions except those dependent on $p$ are linear or quadratic and snap them to be either linear or quadratic manually by using fix_symbolic. After snapping and retraining, the updated KAN is shown in Figure 4.6 (c) step 3, maintaining $97.7\%$ accuracy. From now on, Alice may make two different choices based on her prior knowledge. In one case, Alice may have guessed that the dependence on $p$ is ${\rm cosh}$, so she sets the activations of $p$ to be ${\rm cosh}$ function. She retrains KAN and gets 96.9% accuracy (Figure 4.6 (c) Step 4A). In another case, Alice does not know the ${\rm cosh}\ p$ dependence, so she pursues simplicity and again assumes the functions of $p$ to be quadratic. She retrains KAN and gets 95.4% accuracy (Figure 4.6 (c) Step 4B). If she tried both, she would realize that ${\rm cosh}$ is better in terms of accuracy, while quadratic is better in terms of simplicity. The formulas corresponding to these steps are listed in Table 6. It is clear that the more manual operations are done by Alice, the simpler the symbolic formula is (which slight sacrifice in accuracy). KANs have a “knob" that a user can tune to trade-off between simplicity and accuracy (sometimes simplicity can even lead to better accuracy, as in the GAAM case).

### v3

@@ 2.5.1 Simplification techniques
2. Visualization. When we visualize a KAN, to get a sense of magnitudes, we set the transparency of an activation function $\phi_{l,i,j}$ proportional to ${\rm tanh}(\beta A_{l,i,j})$ where $\beta=3$ . Hence, functions with small magnitude appear faded out to allow us to focus on important ones.
@@
$I_{l,i}=\underset{k}{\rm max}(\left|\phi_{l-1,i,k}\right|_{1}),\qquad O_{l,i}=\underset{j}{\rm max}(\left|\phi_{l+1,j,i}\right|_{1}),$ (2.21)
@@ 3.1 Toy datasets
We train these KANs by increasing grid points every 200 steps, in total covering $G=\{3,5,10,20,50,100,200,500,1000\}$. We train MLPs with different depths and widths as baselines. Both MLPs and KANs are trained with LBFGS for 1800 steps in total. We plot test RMSE as a function of the number of parameters for KANs and MLPs in Figure 3.1, showing that KANs have better scaling curves than MLPs, especially for the high-dimensional example. For comparison, we plot the lines predicted from our KAN theory as red dashed ($\alpha=k+1=4$), and the lines predicted from Sharma & Kaplan [23] as black-dashed ($\alpha=(k+1)/d=4/d$). KANs can almost saturate the steeper red lines, while MLPs struggle to converge even as fast as the slower black lines and plateau quickly. We also note that for the last example, the 2-Layer KAN $[4,9,1]$ behaves much worse than the 3-Layer KAN (shape $[4,2,2,1]$). This highlights the greater expressive power of deeper KANs, which is the same for MLPs: deeper MLPs have more expressive power than shallower ones. Note that we have adopted the vanilla setup where both KANs and MLPs are trained with LBFGS without advanced techniques, e.g., switching between Adam and LBFGS, or boosting [34]. We leave the comparison of KANs and MLPs in advanced setups for future work.
@@ 3.2 Special functions
We collect 15 special functions common in math and physics, summarized in Table 1. We choose MLPs with fixed width 5 or 100 and depths swept in $\{2,3,4,5,6\}$. We run KANs both with and without pruning. KANs without pruning: We fix the shape of KAN, whose width are set to 5 and depths are swept in {2,3,4,5,6}. KAN with pruning. We use the sparsification $(\lambda=10^{-2}\ {\rm or}\ 10^{-3})$ and pruning technique in Section 2.5.1 to obtain a smaller KAN pruned from a fixed-shape KAN. Each KAN is initialized to have $G=3$, trained with LBFGS, with increasing number of grid points every 200 steps to cover $G=\{3,5,10,20,50,100,200\}$. For each hyperparameter combination, we run 3 random seeds.
@@
Name scipy.special API Minimal KAN shape test RMSE $10^{-2}$ Minimal KAN test RMSE Best KAN shape Best KAN test RMSE MLP test RMSE
@@ 3.3 Feynman datasets
I.15.3x $\frac{x-ut}{\sqrt{1-(\frac{u}{c})^{2}}}$ $\frac{1-a}{\sqrt{1-b^{2}}}$ $a,b$ [2,2,1,1] [2,1,1] [2,2,1,1,1] $7.35\times 10^{-3}$ $58\times 10^{-3} 1.1$$14\times 10^{-3}$ $\bf 8.54\times 10^{-4}$}
I.16.6 $\frac{u+v}{1+\frac{uv}{c^{2}}}$ $\frac{a+b}{1+ab}$ $a,b$ [2,2,2,2,2,1] [2,2,1] [2,2,1] $06\times 10^{-3} 1.1$$19\times 10^{-3}$ $1.53\times 10^{-3}$ $\bf 6.20\times 10^{-4}$}
@@
II.2.42 $\frac{k(T_{2}-T_{1})A}{d}$ (a1)b(a-1)b $a,b$ [2,2,1] [2,2,1] [2,2,2,1] $8.54\times 10^{-4}$ $7.22\times 10^{-4}$ $1.22\times 10^{-3}$ $\bf 1.81\times 10^{-4}$}
@@
    KAN with pruning. We use the sparsification $(\lambda=10^{-2}\ {\rm or}\ 10^{-3})$ and the pruning technique from Section 2.5.1 to obtain a smaller KAN from a fixed-shape KAN from (2).
@@
    MLPs with fixed width 5, depths swept in $\{2,3,4,5,6\}$, and activations chosen from $\{{\rm Tanh},{\rm ReLU},{\rm SiLU}\}$.
@@ 3.4 Solving partial differential equations
We compare the KAN architecture with that of MLPs using the same hyperparameters $n_{i}=10000$, $n_{b}=800$, and $\alpha=0.01$. We measure both the error in the $L^{2}$ norm and energy ($H^{1}$) norm and see that KAN achieves a much better scaling law with a smaller error, using smaller networks and fewer parameters; see Figure 3.3. A 2-Layer width-10 KAN is 100 times more accurate than a 4-Layer width-100 MLP ($10^{-7}$ vs $10^{-5}$ MSE) and 100 times more parameter efficient ($10^{2}$ vs $10^{4}$ parameters). Therefore we speculate that KANs might have the potential of serving as a good neural network representation for model reduction of PDEs. However, we want to note that our implementation of KANs are typically 10x slower than MLPs to train. The ground truth being a symbolic formula might be an unfair comparison for MLPs since KANs are good at representing symbolic formulas. In general, KANs and MLPs are good at representing different function classes of PDE solutions, which needs detailed future study to understand their respective boundaries.
@@ 3.5 Continual Learning
Here we simply present our preliminary results on an extremely simple example, to demonstrate how one could possibly leverage locality in KANs (thanks to spline parametrizations) to reduce catastrophic forgetting. However, it remains unclear whether our method can generalize to more realistic setups, especially in high-dimensional cases where it is unclear how to define “locality”. In future work, We would also like to study how our method can be connected to and combined with SOTA methods in continual learning [42, 43].
@@ 4.2 Unsupervised toy dataset
If an algorithm for solving the unsupervised problem can be devised, it has a considerable advantage over the supervised problem, since it requires only the sets of features $S=(x_{1},x_{2},\cdots,x_{d})$. The supervised problem, on the other hand, tries to predict subsets of features in terms of the others, i.e. it splits $S=S_{\text{in}}\cup S_{\text{out}}$ into input and output features of the function to be learned. Without domain expertise to advise the splitting, there are $2^{d}-2$ possibilities such that $|S_{\text{in}}|>0$ and $|S_{\text{out}}|>0$. This exponentially large space of supervised problems can be avoided by using the unsupervised approach. This unsupervised learning approach will be valuable to the knot dataset in Section 4.3. A Google Deepmind team [44] manually chose signature to be the target variable, otherwise they would face this combinatorial problem described above. This raises the question whether we can instead tackle the unsupervised learning directly. We present our method and a toy example below.
We tackle the unsupervised learning problem by turning it into a supervised learning problem on all of the $d$ features, without requiring the choice of a splitting. The essential idea is to learn a function $f(x_{1},\dots,x_{d})=0$ such that $f$ is not the 0-function. To do this, similar to contrastive learning, we define positive samples and negative samples: positive samples are feature vectors of real data. Negative samples are constructed by feature corruption. To ensure that the overall feature distribution for each topological invariant stays the same, we perform feature corruption by random permutation of each feature across the entire training set. Now we want to train a network $g$ such that $g(\mathbf{x}_{\rm real})=1$ and $g(\mathbf{x}_{\rm fake})=0$ which turns the problem into a supervised problem. However, remember that we originally want $f(\mathbf{x}_{\rm real})=0$ and $f(\mathbf{x}_{\rm fake})\neq 0$. We can achieve this by having $g=\sigma\circ f$ where $\sigma(x)={\rm exp}(-\frac{x^{2}}{2w^{2}})$ is a Gaussian function with a small width $w$, which can be conveniently realized by a KAN with shape [,1,1][...,1,1] whose last activation is set to be the Gaussian function $\sigma$ and all previous layers form $f$. Except for the modifications mentioned above, everything else is the same for supervised training.
@@
Now we demonstrate that the unsupervised paradigm works for a synthetic example. Let us consider a 6D dataset, where $(x_{1},x_{2},x_{3})$ are dependent variables such that $x_{3}=\exp(\sin(x_{1})+x_{2}^{2})$; $(x_{4},x_{5})$ are dependent variables with $x_{5}=x_{4}^{3}$; $x_{6}$ is independent of the other variables. In Figure 4.2, we show that for seed = 0, KAN reveals the functional dependence among $x_{1}$$x_{2}$, and $x_{3}$; for another seed = 2024, KAN reveals the functional dependence between $x_{4}$ and $x_{5}$. Our preliminary results rely on randomness (different seeds) to discover different relations; in the future we would like to investigate a more systematic and more controlled way to discover a complete set of relations. Even so, our tool in its current status can provide insights for scientific tasks. We present our results with the knot dataset in Section 4.3.
@@ 4.3 Application to Mathematics: Knot Theory
Knot theory is a subject in low-dimensional topology that sheds light on topological aspects of three-manifolds and four-manifolds and has a variety of applications, including in biology and topological quantum computing. Mathematically, a knot $K$ is an embedding of $S^{1}$ into $S^{3}$. Two knots $K$ and $K^{\prime}$ are topologically equivalent if one can be deformed into the other via deformation of the ambient space $S^{3}$, in which case we write $K]=[K^{\prime}]$. Some knots are topologically trivial, meaning that they can be smoothly deformed to a standard circle. Knots have a variety of deformation-invariant features $f$ called topological invariants, which may be used to show that two knots are topologically inequivalent, $K]\neq[K^{\prime}]$ if $f(K)\neq f(K^{\prime})$. In some cases the topological invariants are geometric in nature. For instance, a hyperbolic knot $K$ has a knot complement $S^{3}\setminus K$ that admits a canonical hyperbolic metric $g$ such that $\text{vol}_{g}(K)$ is a topological invariant known as the hyperbolic volume. Other topological invariants are algebraic in nature, such as the Jones polynomial.
@@
Table 3: KANs can achieve better accuracy than MLPs with much fewer parameters in the signature classification problem. Soon after our preprint was first released, Prof. Shi Lab from Georgia tech discovered that an MLP with only 60 parameters is sufficient to achieve 80% accuracy (public but unpublished results). This is good news for AI + Science because this means perhaps many AI + Science tasks are not that computationally demanding than we might think (either with MLPs or with KANs), hence many new scientific discoveries are possible even on personal laptops.
@@
D $0.09+1.04{\rm exp}(-9.59(-0.62{\rm sin}(0.61\mu_{r}+7.26))-0.32{\rm tan}(0.03\lambda-6.59)+1-0.11e^{-1.77(0.31-\mu_{i})^{2})^{2}}-1.09e^{-7.6(0.65(1-0.01\lambda)^{3}}+0.27{\rm atan}(0.53\mu_{i}-0.6)+0.09+{\rm exp}(-2.58(1-0.36\mu_{r})^{2}))$ $[3,2,1]$ KAN 84.0% 0.947 0.997
@@
    The second group of variables involve cusp volume $V$, real part of meridinal translation $\mu_{r}$ and longitudinal translation $\lambda$. Their activations all look like logarithmic functions (which can be verified by the implied symbolic functionality in Section 2.5.1). So the relation is $\log V+\log\mu_{r}+\log\lambda=0$ which is equivalent to $V=\mu_{r}\lambda$, which is true by definition. It is, however, reassuring that we discover this relation without any prior knowledge.
@@ 4.4 Application to Physics: Anderson localization
KAN auto $\cancel{1.52E^{2}}+21.06\alpha E+\cancel{0.66E}+\cancel{3.55\alpha^{2}}+\cancel{0.91\alpha}+45.13\lambda-54.45=0$ 99.0%
@@
KAN auto $13.99{\rm sin}(0.28{\rm sin}(0.87\lambda+2.22)-0.84{\rm arctan}(0.58E-0.26)+0.85{\rm arctan}(0.94p+0.13)-8.14)-16.74+43.08{\rm exp}(-0.93(0.06(0.13-p)^{2}-0.27{\rm tanh}(0.65E+0.25)+0.63{\rm arctan}(0.54\lambda-0.62)+1)^{2})=0$ 97.1%
KAN man (step 2) + auto $4.19(0.28{\rm sin}(0.97\lambda+2.17)-0.77{\rm arctan}(0.83E-0.19)+{\rm arctan}(0.97p+0.15)-0.35)^{2}-28.93+39.27{\rm exp}(-0.6(0.28{\rm cosh}^{2}(0.49p-0.16)-0.34{\rm arctan}(0.65E+0.51)+0.83{\rm arctan}(0.54\lambda-0.62)+1)^{2})=0$ 97.7%
KAN man (step 3) + auto $4.63E-10.25(-0.94{\rm sin}(0.97\lambda-6.81)+{\rm tanh}(0.8p-0.45)+0.09)^{2}+11.78{\rm sin}(0.76p-1.41)+22.49{\rm arctan}(1.08\lambda-1.32)+31.72=0$ 97.7%
KAN man (step 4A) $6.92E-6.23(-0.92\lambda-1)^{2}+2572.45(-0.05\lambda+0.95{\rm cosh}(0.11p+0.4)-1)^{2}-12.96{\rm cosh}^{2}(0.53p+0.16)+19.89=0$ 96.6%
KAN man (step 4B) $7.25E-8.81(-0.83\lambda-1)^{2}-4.08(-p-0.04)^{2}+12.71(-0.71\lambda+(0.3p+1)^{2}-0.86)^{2}+10.29=0$ 95.4%
@@
Let us assume Alice wants to figure out the mobility edge for MAAM. This task is more complicated and requires more human wisdom. As in the last example, Alice starts from a $[4,2,1,1]$ KAN and trains it but gets an accuracy around 75% which is less than acceptable. She then chooses a larger $[4,3,1,1]$ KAN and successfully gets 98.4% which is acceptable (Figure 4.6 (b) step 1). Alice notices that $\phi$ is not picked up on by KANs, which means that the mobility edge is independent of the phase factor $\phi$ (agreeing with Eq. (4.11)). If Alice turns on the automatic symbolic regression (using a large library consisting of exp, tanh etc.), she would get a complicated formula in Tabel 5-MAAM-KAN auto, which has 97.1% accuracy. However, if Alice wants to find a simpler symbolic formula, she will want to use the manual mode where she does the symbolic snapping by herself. Before that she finds that the $[4,3,1,1]$ KAN after training can then be pruned to be $[4,2,1,1]$, while maintaining $97.7\%$ accuracy (Figure 4.6 (b)). Alice may think that all activation functions except those dependent on $p$ are linear or quadratic and snap them to be either linear or quadratic manually by using fix_symbolic. After snapping and retraining, the updated KAN is shown in Figure 4.6 (c) step 3, maintaining $97.7\%$ accuracy. From now on, Alice may make two different choices based on her prior knowledge. In one case, Alice may have guessed that the dependence on $p$ is $\rm cosh$}, so she sets the activations of $p$ to be $\rm cosh$} function. She retrains KAN and gets 96.9% accuracy (Figure 4.6 (c) Step 4A). In another case, Alice does not know the ${\rm cosh}\ p$ dependence, so she pursues simplicity and again assumes the functions of $p$ to be quadratic. She retrains KAN and gets 95.4% accuracy (Figure 4.6 (c) Step 4B). If she tried both, she would realize that $\rm cosh$} is better in terms of accuracy, while quadratic is better in terms of simplicity. The formulas corresponding to these steps are listed in Table 5. It is clear that the more manual operations are done by Alice, the simpler the symbolic formula is (which slight sacrifice in accuracy). KANs have a “knob" that a user can tune to trade-off between simplicity and accuracy (sometimes simplicity can even lead to better accuracy, as in the GAAM case).